OpenStudy (anonymous):

how do you solve x+1/3-8/3 ?

OpenStudy (anonymous):

do you have what the equation equals?

OpenStudy (anonymous):

i'm assuming you trying to find x

OpenStudy (anonymous):

yes I'm expected to solve for x

OpenStudy (anonymous):

did the question not tell you what that whole thing equals?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[\frac{ x+1 }{ 3 }-\frac{ 8 }{ 3 }= y\]

OpenStudy (anonymous):

\[\frac{ x+1 }{ 3 } = \frac{ -8 }{ 3 }\]

OpenStudy (anonymous):

x+1 = -8

OpenStudy (anonymous):

x = -9

OpenStudy (anonymous):

alright Thank you! I understand now

OpenStudy (anonymous):

yw :)